# Channel Balance Interpolation in the Lightning Network via Machine Learning thanks: * Jointly supervised

### v1

@@ VIII-A1 Equal Split
Split the capacity equally in the two channels:
@@ VIII-B1 Edge-Wise Random Features Prediction
Assign random features $r_{(u,v)}$ from an isotropic Normal distribution to each edge. Train a model which predicts $p_{(u,v)}$ only as a function of the random features of the edge $(u,v)$:
@@ VIII-B2 Node-Wise Prediction
A model which predicts $p_{(u,v)}$ only as a function of the features of source node $u$:
@@ VIII-C Methodology
  • $MAE_{y}$: Mean Absolute Error between the actual edge balances and the predictions.
  • $MAE_{p}$: Mean Absolute Error in the predictions of edge balances.
  • $R$: The correlation coefficient, indicating how closely the predictions align with the actual values.
  • $R^{2}$: The coefficient of determination, showing the percentage of the variance in the actual edge balances that is explained by the model.
The model could be any kind of machine learning model. In this study, we used a Random Forest regression model.
This table allows for a straightforward comparison of model performance, focusing on accuracy (MAE) and the strength of the relationship between predicted and actual values (R and R²). The goal being to identify the model that offers the best balance between minimizing prediction errors and maximizing explanatory power for edge balances.
Additionally, the Mean Decrease Impurity (MDI) feature importance was plotted, illustrating the contribution of each feature to model effectiveness.
@@ IX Results
The results are recorded in a table in Figure 5. Note that MAEy is denoted in millions of satoshis.
@@
Random Edge Features $0.362$ $1.67$ $-0.06$ $-0.03$
@@
Figure 5: Performance summary of all models and heuristics.
The heuristics Equal Split and Local Max HTLC provide initial reference points for comparing performance. The Equal Split and Local Max HTLC models demonstrate negligible predictive power, explaining virtually none of the variance in the dependent variable. The Node-wise and Edge-wise models show some predictive ability but still limited effectiveness. Moderate improvements are seen in the Concatenated and Shallow Graph models. The Joint model, which considers node, edge, and positional encoding features, has the greatest predictive power.
The inclusion of positional encodings in the Random Forest model resulted in improved performance compared to the models without positional encodings. This suggests that positional encodings enhance the model’s ability to capture the underlying data structure. Additionally, it is noteworthy that the Shallow Graph model outperformed the Equal Split heuristic, indicating that positional information alone is predictive of channel balances.
@@
The histogram in Figure 6 presents the error distribution for the top-performing model from the analysis. Errors are binned between -1 and 1, and the y-axis is the frequency of each errors. The model’s errors presents a distribution that peaks around the central value, zero. As the histogram tapers off on both sides, so does the frequency of larger errors. The symmetry of the data indicates that the model does not consistently overpredict or underpredict.
The scatterplot in Figure 6 presents a comparison between actual and predicted balance percentages. The best fit line for this scatterplot has a positive slope indicating that the predicted values tend to increase as the actual values increase, although this is clearly not a tight correlation.
From the confusion matrix in Figure 6, it is apparent that the model is able to identify which side a majority of the liquidity is present. For example, of the channels in the test data that have less local liquidity than remote, the model predicted a $\hat{p}\leq 0.5$ 72% of the time.
@@ IX-A Feature Importance
Notably, Positional Encoding emerges as the most influential feature, suggesting that the model heavily weighs the structural positioning of nodes within the network. Other network features like Capacity Centrality highlight the strategic importance of controlling capacity within the network’s topology.
While the economic features, Fee Rate, Remote Fee, Fee Ratio, and Local Fee have lower MDI values compared to the network features, they still depict how the cost dynamics affect the model’s decision-making process. The feature Max HTLC (%), despite its lower ranking, indicates that constraint on payment size is useful in predicting channel balances. These economic features reveal how the fee structures and transactional ceilings set by the channels incentivize certain flow patterns and impact the distribution of funds across the network.
The economic features and network features each hold approximately half of the importance. This suggests a balanced reliance on both economic and network attributes. Together, these network and economic features enhance the model’s ability to provide a nuanced prediction of channel balances.
@@ X Discussion
This study was conducted to explore the utility of machine learning models in predicting channel balances within the Lightning Network. A novel aspect to this method is the incorporation of graph features into the machine learning models. This approach leverages the inherent structure of the network, utilizing the relational and topological data that characterizes the connections between nodes and channels. The results suggest that machine learning techniques, such as regression algorithms and neural networks, can indeed provide accurate predictions of channel balances. This indicates a promising direction for further research and potential practical application in enhancing the efficiency of network operations.
Probing techniques to estimate channel balances, while useful, often incur significant overhead and can potentially compromise user privacy due to the frequent and sometimes intrusive nature of probes. In contrast, our balance interpolation approach uses only crowdsourced data and reduces the need for frequent probing by providing a probabilistic estimation of balances that can be updated as new transaction data becomes available. This method not only diminishes the operational overhead associated with probing but also enhances user privacy by reducing the frequency of direct channel queries.
@@ XI Future Work
A natural follow up to this experiment is to implement an enhanced pathfinding algorithm that integrates the predictive model into a Lightning Node’s routing decisions. The Enhanced Pathfinding Algorithm aims to find the most reliable path from a source node to a destination node. It utilizes a three-step process: first, predict the balance of each edge in the network; second, assign a cost to each edge according to its predicted balance; and finally, use Djikstras algorithm to find the cheapest path. The cost function used is the negative logarithm which has been shown in other works to be useful for optimizing pathfinding for reliability[3].
Algorithm 1 Enhanced Pathfinding Using Predictive Model
0:  $G(V,E)$, a graph where $V$ are nodes and $E$ are edges.
0:  $src$, the source node.
0:  $dest$, the destination node.
0:  $amt$, the payment amount.
0:  M, a model that predicts the balance of each channel.
0:  The most reliable path from $s$ to $d$ based on model predictions.
1:  $\mathcal{C}\leftarrow$ empty map {Store the cost for each edge}
2:  for all edge $(u,v)\in E$ do
3:     $feats\leftarrow$ ExtractFeatures($(u,v)$)
4:     $\hat{p}\leftarrow$ M.predict($(u,v)$,$feats$)
5:     if  EffectiveBalance($(u,v)$, $\hat{p}$) $<amt$ then
6:        $\mathcal{C}[(u,v)]\leftarrow\infty$
7:     else
8:        $\mathcal{C}[(u,v)]\leftarrow-\log(\hat{p})$Convert balance to cost}
9:     end if
10:  end for
11:  $path\leftarrow Dijkstra({G,src,dest,\mathcal{C}})$
12:  return  $path$
Another future direction of work is to conduct analysis via simulation to determine how the number of payment retries changes when using this machine learning model versus existing pathfinding methods. This can be directly translated to time saved in waiting for payment confirmation, a critical aspect of Lightning wallet user experience (UX). This quantified pathfinding efficiency gain is expected to go up as interpolation models improve.
By adding more features or more diverse historical data, the model can potentially uncover more intricate patterns and improve its predictions. For example, adding data related to transaction frequencies, channel opening and closing times, or network congestion might provide deeper insights into balance dynamics. Incorporating time series analysis could help capture trends and seasonal variations in channel balances. This approach could make predictions more dynamic by considering how balances change, not just the static state of balances at a single point in time.
In addition to supervised learning, the probabilistic learning of distributions would be useful for conducting simulations so balances can be repeatedly sampled to generate robust confidence intervals. Robust confidence intervals, derived from well-modeled probabilistic distributions, can give network operators and users more reliable information for path planning. By addressing these areas, the model not only becomes more accurate in its predictions but also more useful for planning and operational decisions within the Lightning Network. This enhancement could lead to more efficient routing decisions, reduced transaction costs, and improved overall network stability.

### v2

@@ VIII Model Evaluation
Our primary objective is to minimize the prediction error of channel balances while ensuring the model generalizes well across different network conditions. Specifically, we seek to learn a function $f_{\Theta}$ that predicts the balance proportion $\hat{p}_{(u,v)}$ for each edge $(u,v)$ using node, edge and network features.
@@ VIII-A1 Equal Split
Split the capacity equally between the two edges:
@@ VIII-B1 Edge-Wise Random Features Prediction
Assign random features $r_{(u,v)}$ from an isotropic Normal distribution to each edge. Train a model that predicts $p_{(u,v)}$ only as a function of the random features of the edge $(u,v)$:
@@ VIII-B2 Node-Wise Prediction
A model which predicts $p_{(u,v)}$ only as a function of the features of the source node $u$:
@@ VIII-B5 Shallow Graph Prediction
VIII-B6 Joint Prediction
A model which predicts $p_{(u,v)}$ as a function of the concatenation of the node features, edge features and the positional encodings:
[…]
@@ VIII-C Methodology
  • $MAE_{y}$: Mean Absolute Error between actual and predicted edge balances.
  • $MAE_{p}$: Mean Absolute Error in edge balance predictions.
  • $R$: Correlation coefficient, measuring the alignment between predictions and actual values.
  • $R^{2}$: Coefficient of determination, representing the proportion of variance in edge balances explained by the model.
This table allows for a straightforward comparison of model performance, focusing on accuracy ($MAE$) and the strength of the relationship between predicted and actual values ($R$ and $R^{2}$). The goal being to identify the model that offers the best balance between minimizing prediction errors and maximizing explanatory power for edge balances. Additionally, the Mean Decrease Impurity (MDI) feature importance was plotted, highlighting each feature’s contribution to the model’s effectiveness.
@@ IX Results
The results are recorded in Table II. Note that MAEy is denoted in millions of satoshis.
TABLE II: Model Performance
@@
Random Edge Features $0.362$ $1.67$ 0.06-0.06 0.03-0.03
@@
The Equal Split and Local Max HTLC heuristics provide initial reference points for comparing performance. The Equal Split and Local Max HTLC models demonstrate negligible predictive power, explaining virtually none of the variance in the dependent variable. The Node-wise and Edge-wise models show some predictive ability but still limited effectiveness. Moderate improvements are seen in the Concatenated and Shallow Graph models. The Joint model, which considers node, edge, and positional encoding features, has the greatest predictive power.
The inclusion of positional encodings in the Random Forest model resulted in improved performance compared to models without positional encodings. This suggests that positional encodings enhance the model’s ability to capture the underlying data structure. Additionally, it should be noted that the shallow graph model outperformed the Equal Split heuristic, indicating that positional information alone is predictive of channel balances.
@@
The histogram in Figure 4 illustrates the error distribution for the top-performing model. Errors are binned between -1 and 1, with the y-axis representing their frequency. The distribution peaks around zero, tapering off symmetrically on both sides, indicating that large errors are less frequent and that the model does not systematically overpredict or underpredict.
The scatterplot in Figure 4 compares actual and predicted balance percentages. The best-fit line has a positive slope, suggesting that predicted values generally increase with actual values, though the correlation is not particularly strong.
The confusion matrix in Figure 4 reveals that the model effectively identifies which side holds the majority of the liquidity. Notably, for test data where there is less local liquidity than remote, the model correctly predicted $\hat{p}\leq 0.5$ 72% of the time. The resulting precision is $0.74$ and the recall is $0.7255$.
@@ IX-A Feature Importance
TABLE III: Feature importance using mean decrease in impurity
Feature Importance (%)
Max HTLC 5.0
Local Fee 8.3
Remote Fee 12.5
Fee Ratio 10.2
Fee Rate (ppm) 14.7
Capacity Centrality 16.0
Positional Encoding 33.3
Notably, Positional Encoding emerges as the most influential feature, indicating that the model heavily relies on the structural positioning of nodes within the network. Other network-related features, such as Capacity Centrality, underscore the predictive significance of controlling capacity within the network’s topology.
While economic features Fee Rate, Remote Fee, Fee Ratio, and Local Fee exhibit lower MDI values compared to network-based features, they still demonstrate the role of cost dynamics in shaping the model’s decision-making process. Max HTLC (%), despite ranking lower, suggests that constraints on payment size contribute to predicting channel balances. These economic features highlight how fee structures and transactional limits influence fund distribution and flow patterns within the network.
Overall, economic and network features contribute nearly equally to model importance, suggesting a balanced reliance on both structural and financial attributes. Together, these factors enhance the model’s ability to make nuanced predictions of channel balances.
@@ X Discussion
This study was conducted to explore the utility of machine learning models in predicting channel balances within the Lightning Network. A novel aspect to this method is the incorporation of graph features into the machine learning models. The results suggest that machine learning techniques can provide accurate predictions of channel balances.
While probing techniques can estimate channel balances, they often introduce significant overhead and pose potential privacy risks. In contrast, our balance interpolation approach relies solely on crowdsourced data, offering a probabilistic estimation that minimizes the need for frequent probing. This reduces operational overhead and enhances user privacy by decreasing direct channel queries.
@@ XI Future Work
A natural next step is to develop an enhanced pathfinding algorithm that integrates the predictive model into a Lightning Node’s routing decisions. The enhanced pathfinding algorithm would identify the most reliable path from a source to a destination through a three-step process:
  1. 1.
    Predict the balance of each edge in the network.
  2. 2.
    Assign a cost to each edge based on its predicted balance.
  3. 3.
    Compute the optimal path using Dijkstra’s algorithm, selecting the path with the lowest cost.
The cost function is the negative logarithm of the predicted balance, a technique shown in prior work to be effective for optimizing pathfinding reliability even with uncertain channel balances[3].